# EmoDynamiX: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics

### v1

@@ Abstract
Designing emotionally intelligent conversational systems to provide comfort and advice to people experiencing distress is a compelling area of research. Previous efforts have focused on developing modular dialogue systems that treat socio-emotional strategy prediction as an auxiliary task and generate strategy-conditioned responses with customized decoders. Recently, with advancements in large language models (LLMs), end-to-end dialogue agents without explicit socio-emotional strategy prediction steps have become prevalent. However, despite their excellence in language generation, recent studies show that LLMs’ inherent preference bias towards certain socio-emotional strategies hinders the delivery of high-quality emotional support. To address this challenge, we propose decoupling strategy prediction from language generation, and introduce a novel dialogue strategy predictor, EmoDynamiX, which models the discourse dynamics between user emotions and system strategies using a heterogeneous graph. Additionally, we make use of the Emotion Recognition in Conversations (ERC) task and design a flexible mixed-emotion module to capture fine-grained emotional states of the user. Experimental results on two ESC datasets show EmoDynamiX outperforms previous state-of-the-art methods with a significant margin. Our code is available at https://github.com/cw-wan/EmoDynamiX-v2.
@@ 1 Introduction
Figure 1: An example of an emotional support conversation between the user and the conversational agent. Our task is to predict the next socio-emotional strategy for the agent.
@@
Previous work on data-driven ESC has primarily focused on modular dialogue systems as defined by Clavel et al. (2022). Modular systems feature a three-fold workflow: recognizing, planning, and generating. Examples include Tu et al. (2022), Deng et al. (2023), Liu et al. (2021) and Cheng et al. (2022). These systems select the socio-emotional strategy based on the user’s state and generate responses conditioned on the predicted strategy. In this paradigm, a strategy planner is usually optimized with a customized language decoder in either a parallel or sequential manner. However, following the advent of Large Language Models (LLMs) with stronger generative abilities, end-to-end conversational systems have started to garner more attention (Chen et al., 2023b; Zheng et al., 2023), which implies that strategy planning transitions from an explicit process to a hidden mechanism. Despite this shift, recent studies indicate that LLMs are not sophisticated socio-emotional strategy planners, as their inherent bias towards certain strategies prevents them from balancing social-oriented and task-oriented goals. For example, a study by Abulimiti et al. (2023) shows that LLMs tend to push forward the dialogue process when strategies to amend rapport should be applied, resulting in poor conversation outcomes.
Clues for mitigating the limitations of LLM-based end-to-end dialogue systems may lie in modular dialogue systems that explicitly predict strategies. Studies across various dialogue tasks demonstrate a positive correlation between high-quality socio-emotional strategy prediction and the successful fulfillment of conversational tasks (Liu et al., 2021; He et al., 2018; Joshi et al., 2021; Raut et al., 2023). Therefore, we propose treating socio-emotional strategy prediction as an independent task (Figure 1) as previously done in Vanel et al. (2023). Our aim is to develop high-quality strategy predictors that could benefit LLM-based conversation agents in the long run.
We propose EmoDynamiX, a novel socio-emotional strategy prediction framework. EmoDynamiX introduces two key innovations: fine-grained user emotion state modelling and the use of heterogeneous graphs to capture conversational dynamics. Firstly, we make use of the Emotion Recognition in Conversations (ERC) task by designing a flexible mixed-emotion module that transfers knowledge from emotion distributions predicted by a pre-trained ERC model. Additionally, we integrate user emotional states and previous agent strategies into a heterogeneous graph, effectively capturing the transitions of user emotions and their interactions with past strategies. While graph learning is widely applied in conversational recognition tasks, it remains relatively underexplored in predictive tasks. To address this, we utilize dummy nodes (Liu et al., 2022; Scarselli et al., 2008) for role-aware information aggregation, enhancing the overall performance.
Our contributions are summarized as follows: 1) We propose a socio-emotional strategy prediction framework that effectively models conversational dynamics with fine-grained emotional states. 2) We introduce role-aware dummy nodes to improve the performance of graph-learning in our predictive task. 3) Experimental results on two public ESC datasets demonstrate that EmoDynamiX outperforms all previous baselines by a significant margin, achieving over 10% improvement on ESConv (Liu et al., 2021) and over 30% improvement on the low-resource dataset AnnoMI (Wu et al., 2022).
@@ 2.1 Emotional Support Conversation
Emotional Support Conversation aims to alleviate distress by expressing empathy and providing suggestions. The goal is both social-oriented and task-oriented. Liu et al. (2021) introduced a joint-learning baseline to condition response generation with predicted strategies. Cheng et al. (2023) infer the user’s persona from dialogue history to create more personalized ESC systems. Cheng et al. (2022) and Li et al. (2024) focus on identifying strategies that optimize future user feedback.
Modelling user states is a critical topic in ESC. Previous work commonly approaches this by querying commonsense knowledge graphs (Tu et al., 2022; Deng et al., 2023; Peng et al., 2022; Zhao et al., 2023; Li et al., 2024). These queries are constructed by concatenating the current utterance with specific knowledge relations, such as xReact and xEffect. The queries are then fed into COMET (Hwang et al., 2021), a generative model pre-trained on commonsense knowledge graphs, which returns the user’s reaction (xReact) or feeling (xEffect) to the situation described in the utterance. However, commonsense knowledge is too general to capture the fine-grained emotional states entailed in specific speaker turns like "Okay". In contrast, models specialized in dialogue with context-aware architectures, such as sequential or graph-based models, trained on Emotion Recognition in Conversation (ERC) datasets, can handle these nuances more effectively. Additionally, emotions are frequently mixed in real-life situations, and contradictory emotions (like Sadness and Joy) could coexist in specific contexts (Braniecka et al., 2014). Our mixed-emotion modelling approach handles this complexity better and can model a large set of subtle emotional expressions by combining primary emotions without further human annotations (as demonstrated in Section 5.4).
@@ 2.2 Graph Learning in Conversational Tasks
Graph-based approaches have proven effective in various dialogue-related tasks. In recognition tasks, such as conversational emotion recognition and dialogue act recognition, the target speaker turn aggregates information from its neighbors according to the graph structure. Studies by Ghosal et al. (2019), Ishiwatari et al. (2020), Wang et al. (2020), Fu et al. (2023), and Shen et al. (2021) design dialogue graphs based on interactions between speaker roles. Li et al. (2023) and Zhang et al. (2023) construct dialogue graphs based on discourse dependencies parsed with a pre-trained expert model, an approach also applied by Chen and Yang (2021) and Feng et al. (2021) in dialogue summarization. Yang et al. (2023) incorporate commonsense knowledge as heterogeneous nodes. Furthermore, Hu et al. (2021) and Chen et al. (2023a) model multi-modal fusion in dialogue graphs.
In predictive dialogue tasks, such as predicting the next dialogue strategy or act, predictions must rely on global information extracted from graphs. Previous works have utilized simple readout functions, such as mean/max pooling (Joshi et al., 2021) and linear layers (Raut et al., 2023). Our approach innovates by introducing dummy nodes as special placeholders for information aggregation. While dummy nodes have been previously used in other graph-learning tasks, such as graph classification and subgraph isomorphism matching (Liu et al., 2022), they have primarily served as alternatives to readout functions, which are used for obtaining embeddings of graphs or subgraphs. We are the first to employ dummy nodes in a predictive dialogue task, which is particularly useful, since it allows to clearly model role-aware interactions with previous speaker turns.
@@ 3 Problem Formulation
The task of predicting the next dialogue strategy can be written as a multi-class classification problem: assuming a dialogue comprising $T$ speaker turns, we define the dialogue history as $H^{T}=\{U^{T},A^{T},ST^{T}\}$ where $U^{T}=\{u_{t}\}_{t=1}^{T}$ is the sequence of utterances, and each $u_{t}=\{w_{n}\}_{n=1}^{N^{t}}$ is a sequence of $N^{t}$ words. $A^{T}=\{a_{t}\}_{t=1}^{T}$is the sequence of speaker roles, with $a_{t}\in\{\texttt{user},\texttt{system}\}$. $ST^{T}=\{{st}_{t}\}_{t=1}^{T}$ is the sequence of possible strategies, but they only exist for the agent. Noting $\mathcal{S}$ the set of strategies, $I_{\texttt{user}}=\left\{t,a_{t}=\texttt{user}\right\}$ the indexes of turns where the user is speaking, and denoting similarly $I_{\texttt{agent}}$, we have that $\forall t\in I_{\texttt{agent}},{st}_{t}\in\mathcal{S}$ and $\forall t\in I_{\texttt{user}},{st}_{t}=\varnothing$. Our task is, given a fixed window size of $N-1$, to predict the strategy for agent speaker turns, which we formulate as estimating the probability distribution $\mathbb{P}(st_{t}\mid H_{t-N+1}^{t-1})$ upon $\mathcal{S}$ when $t\in I_{\texttt{agent}}$. In the next section, in order to keep our notations simple, we describe our model with indexes corresponding to estimating $\mathbb{P}(st_{N}\mid H_{1}^{N-1})$.
@@ 4 Methodology
The overview of our proposed framework is illustrated in Figure 2. Our framework comprises three main components: (1) a semantic modelling module for capturing the global semantic feature in the dialogue context; (2) a heterogeneous graph learning module, designed to capture the complex interplay between user emotions and system strategies within the dialogue history; and (3) an information fusion module, which integrates the features obtained from the previous modules to produce the prediction result.
Figure 2: The overview of our proposed model which consists of a semantic modelling module, a heterogeneous graph learning module and an information fusion module.
@@ 4.2 Heterogeneous Graph Learning
We propose to use a heterogeneous graph (HG) to model the dynamic interactions contained in the history $H_{1}^{N-1}$. This graph is formally defined as $\mathcal{G}=\{\mathcal{V},\mathcal{B}\}$, which are respectively the set of nodes and edges. Both can be of several types: broadly, node types correspond to past strategies adopted by the conversational agent, the user’s previous emotional states, and the strategy to be predicted; while edge types model the discourse dependencies between dialogue turns and facilitate information aggregation.
Node Types: Our heterogeneous graph contains $N$ nodes $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$, each corresponding to a speaker turn, including the target turn $N$. For user turns $i\in I_{\texttt{user}}$, $v_{i}$ is an emotion node, which encapsulate the fine-grained emotion state of the user. For user turns $i\in I_{\texttt{agent}}$, $v_{i}$ is a system strategy node, that represent the specific conversational strategy implemented by the agent. Lastly, we introduce the dummy node $v_{N}$ as a placeholder for the target utterance, aggregating information from the two other node types and their interactions.
Edge Types: The edges in our heterogeneous graph fulfill dual roles: they model the discourse dependencies between dialogue turns and facilitate information aggregation towards the dummy node. Discourse dependencies correspond to edges between and within emotion and system strategy nodes: we follow Asher et al. (2016)’s definition (which include categories such as Comment and Elaboration) and pre-train a discourse parser as proposed by Chi and Rudnicky (2022) on the multi-party discourse dataset STAC (Asher et al., 2016). Details regarding the training procedure can be found in Appendix B. We note $\mathcal{R}_{\text{Discourse}}$ the set of possible dependencies; then, $\forall(i,j)$ such that $1\leq i,j\leq N-1,\langle v_{i},v_{j}\rangle\in\mathcal{R}_{\text{Discourse}}$. We give more details on these dependencies in Appendix A. The remaining edges are aggregating information from system strategy nodes, which we call self-reference: $\forall i\in I_{\texttt{agent}},\langle v_{i},v_{N}\rangle=r_{\text{self}}$; and from user emotion nodes, which we call inter-reference: $\forall i\in I_{\texttt{user}},\langle v_{i},v_{N}\rangle=r_{\text{inter}}$. We note $\mathcal{R}=\mathcal{R}_{\text{Discourse}}\cup\{r_{\text{self}},r_{\text{inter}}\}$ the set of edge types.
@@ 4.2.1 User Emotion Node Embedding: Mixed Emotion Method
Unlike previous work that relies on commonsense knowledge, we propose leveraging a pre-trained Emotion Recognition in Conversation (ERC) model to predict emotion distributions from user utterances. We then utilize the knowledge contained in these distributions to create embeddings for fine-grained user states using a mixed-prototype approach.
@@
For DailyDialog, we categorize emotions into seven groups: Ekman’s six basic emotions plus Neutral, collectively referred to as $\mathcal{E}$. Additional statistical details about DailyDialog, along with our motivation for selecting it, are available in Appendix A. Detailed information regarding the implementation and training hyperparameters can be found in Appendix B.
@@
To utilize the information in the emotion label distribution more effectively, we employ a learnable temperature parameter $\tau$; details on the impact of the initialization of $\tau$ can be found in Appendix G.

### v2

@@ Abstract
Designing emotionally intelligent conversational systems to provide comfort and advice to people experiencing distress is a compelling area of research. Recently, with advancements in large language models (LLMs), end-to-end dialogue agents without explicit strategy prediction steps have become prevalent. However, implicit strategy planning lacks transparency, and recent studies show that LLMs’ inherent preference bias towards certain socio-emotional strategies hinders the delivery of high-quality emotional support. To address this challenge, we propose decoupling strategy prediction from language generation, and introduce a novel dialogue strategy prediction framework, EmoDynamiX, which models the discourse dynamics between user fine-grained emotions and system strategies using a heterogeneous graph for better performance and transparency [^1] Our code is available at https://anonymous.4open.science/r/EmoDynamiX-v2-F6C6.. Experimental results on two ESC datasets show EmoDynamiX outperforms previous state-of-the-art methods with a significant margin (better proficiency and lower preference bias). Our approach also exhibits better transparency by allowing backtracing of decision making.
@@ 1 Introduction
Figure 1: This figure shows a possible modular ESC dialogue system, in which our dialogue strategy prediction framework outputs the next dialogue strategy to be used to guide an external generative model.
@@
Previous work on data-driven ESC has primarily focused on modular dialogue systems as defined by Clavel et al. (2022). Modular systems feature a three-fold workflow: recognizing, planning, and generating. Examples include Tu et al. (2022), Deng et al. (2023), Liu et al. (2021) and Cheng et al. (2022). In these systems, the socio-emotional strategy is selected based on the recognition of the user’s state, and responses are generated using customized language decoders conditioned on the predicted strategy. With the emergence of Large Language Models (LLMs) offering enhanced capabilities, LLMs have increasingly dominated both modular and, particularly, end-to-end ESC systems (Chen et al., 2023b; Zheng et al., 2023), where strategy planning has shifted from an explicit process to a more implicit, hidden mechanism.
However, implicit dialogue strategy planning with LLMs faces two challenges. First, transparency is often lacking in such "black-box" decision-making processes. Second, recent studies show that preference biases inherited from pre-training data often cause LLMs to struggle with balancing social-oriented and task-oriented goals. Abulimiti et al. (2023) found that in peer-tutoring dialogues, LLMs like ChatGPT frequently prioritize non-hedging strategies, even in situations where hedging strategies would be more appropriate for repairing low rapport between peers. Similarly, Kang et al. (2024) observed that LLMs’ strong predisposition towards certain strategies can undermine the outcome of the current stage of ESC. As a result, the overall objectives may be significantly compromised.
To address this limitation, introducing external strategy planners, which offer greater controllability by enabling us to explicitly exclude inappropriate strategies in specific contexts, stands out to be a promising solution. It has been evidenced by both automatic metrics and human evaluations that, an external decision agent can more effectively mitigate preference bias and improve emotional support quality compared to relying solely on enhancing end-to-end LLMs (Kang et al., 2024). Therefore, we treat socio-emotional strategy prediction as an independent task, as previously explored by Vanel et al. (2023). We illustrate the scope of this task in Figure 1. From this perspective, we raise the following research questions:
  • RQ1: Can we better enhance performance and transparency in socio-emotional dialogue strategy prediction of modular systems by smartly designing dedicated frameworks, rather than prompting or fine-tuning LLMs?
  • RQ2: Given the importance of emotional intelligence in delivering effective emotional support, can we boost strategy prediction in ESC by integrating the task of Emotion Recognition in Conversations (ERC)?
In addressing RQ1, we introduce EmoDynamiX, a decision agent that integrates multiple expert models and incorporates a heterogeneous graph learning module to capture the dynamic interactions between system strategies and user emotions. With graphs, we backtrace the decision-making process, making a step towards greater transparency. We also utilize dummy nodes (Liu et al., 2022; Scarselli et al., 2008) for role-aware information aggregation, enhancing the overall performance.
For RQ2, we design a mixed-emotion module to effectively integrate ERC into our framework: (1) By using emotion distributions instead of discrete labels, we reduce the risk of error propagation, as there are domain gaps between ERC and ESC datasets. (2) By tuning emotion distributions, we can effectively model nuanced emotion categories by fusing primary emotions.
We validate the effectiveness of our proposed framework through comparative experiments on two public ESC datasets. The results demonstrate that EmoDynamiX significantly outperforms all previous baselines, achieving superior F1 scores and a notable reduction in preference bias.
@@ 2.1 Emotional Support Conversation
The goal of ESC is both social-oriented and task-oriented. It aims to alleviate distress by expressing empathy and providing suggestions Liu et al. (2021); Cheng et al. (2022). Modelling user states is thus a critical topic in ESC. Previous work commonly approaches this by querying commonsense knowledge graphs (Tu et al., 2022; Deng et al., 2023; Peng et al., 2022; Zhao et al., 2023; Li et al., 2024). These queries are constructed by concatenating the current utterance with specific knowledge relations, such as xReact. The queries are then fed into COMET (Hwang et al., 2021), a generative model pre-trained on commonsense knowledge graphs, which returns the user’s emotional reaction (xReact) to the situation. However, commonsense knowledge is too general to capture the fine-grained emotional states.
In contrast, models specialized in dialogue with context-aware architectures, such as sequential or graph-based models, trained on ERC datasets, can handle these nuances more effectively. Additionally, emotions are frequently mixed in real-life situations, and contradictory emotions (like Sadness and Joy) could coexist in specific contexts (Braniecka et al., 2014). Our mixed-emotion modelling approach handles this complexity better and can model a large set of subtle emotional expressions by combining primary emotions without further human annotations (as demonstrated in Section 6).
@@ 2.2 Graph Learning in Conversational Tasks
Graph-based approaches have proven effective in various dialogue-related tasks. In recognition tasks, such as conversational emotion recognition and dialogue act recognition, the target speaker turn aggregates information from its neighbors according to the graph structure. Studies by Ghosal et al. (2019), Ishiwatari et al. (2020), Wang et al. (2020), Fu et al. (2023), and Shen et al. (2021) design dialogue graphs based on interactions between speaker roles. Li et al. (2023) and Zhang et al. (2023) construct dialogue graphs based on discourse dependencies parsed with a pre-trained expert model, an approach also applied by Chen and Yang (2021) and Feng et al. (2021) in dialogue summarization. Yang et al. (2023) incorporate commonsense knowledge as heterogeneous nodes. Furthermore, Hu et al. (2021b) and Chen et al. (2023a) model multi-modal fusion in dialogue graphs.
In predictive dialogue tasks, such as forcasting the next dialogue act, decisions rely on global information extracted from graphs. Previous works have utilized simple readout functions, such as mean/max pooling (Joshi et al., 2021) and linear layers (Raut et al., 2023). Our approach introduces dummy nodes as special placeholders for information aggregation. While dummy nodes have been previously used in other graph-learning tasks, such as graph classification and subgraph isomorphism matching (Liu et al., 2022), they have primarily served as alternatives to readout functions, which are used for obtaining embeddings of graphs or subgraphs. We are the first to employ dummy nodes in a predictive dialogue task, which is particularly useful, since it allows to clearly model role-aware interactions with previous speaker turns.
@@ 3 Problem Formulation
The task of predicting the next dialogue strategy can be written as a multi-class classification problem: assuming a dialogue comprising $T$ speaker turns, we define the dialogue history as $H^{T}=\{U^{T},A^{T},ST^{T}\}$ where $U^{T}=\{u_{t}\}_{t=1}^{T}$ is the sequence of utterances, and each $u_{t}=\{w_{n}\}_{n=1}^{N^{t}}$ is a sequence of $N^{t}$ words. $A^{T}=\{a_{t}\}_{t=1}^{T}$is the sequence of speaker roles, with $a_{t}\in\{\texttt{user},\texttt{system}\}$. $ST^{T}=\{{st}_{t}\}_{t=1}^{T}$ is the sequence of possible strategies, but they only exist for the agent. Noting $\mathcal{S}$ the set of strategies, $I_{\texttt{user}}=\left\{t,a_{t}=\texttt{user}\right\}$ the indexes of turns where the user is speaking, and denoting similarly $I_{\texttt{agent}}$, we have that $\forall t\in I_{\texttt{agent}},{st}_{t}\in\mathcal{S}$ and $\forall t\in I_{\texttt{user}},{st}_{t}=\varnothing$. Our task is, given a fixed window size of $N-1$, to predict the strategy for agent speaker turns, which we formulate as estimating the probability distribution $\mathbb{P}(st_{N}\mid H_{1}^{N-1})$ upon $\mathcal{S}$ when $t\in I_{\texttt{agent}}$. For the sake of simplicity, from now on we will use indexing from the beginning of the context window and no longer from the entire conversation.
@@ 4 Methodology
Our framework (see Figure 2) comprises three main components: (1) a semantic modelling module for capturing the semantics of the dialogue context; (2) a heterogeneous graph learning module, designed to capture the complex interplay between the user’s emotions and system strategies within the dialogue history; and (3) an MLP classification head which integrates the features obtained from the previous modules to produce the prediction result.
Figure 2: The overview of our proposed model that consists of a semantic modelling module, a heterogeneous graph learning module, and an MLP classification head.
@@ 4.2 Heterogeneous Graph Learning
We propose to use a heterogeneous graph (HG) to model the interaction dynamics in the history $H_{1}^{N-1}$. This graph is formally defined as $\mathcal{G}=\{\mathcal{V},\mathcal{B}\}$, which are respectively the set of nodes and edges. Both can be of several types: broadly, node types correspond to past strategies adopted by the conversational agent, the user’s previous emotional states, and the strategy to be predicted; while edge types model the discourse dependencies between dialogue turns and facilitate information aggregation.
Node Types: Our heterogeneous graph contains $N$ nodes $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$, each corresponding to a speaker turn, including the target turn $N$. For user turns $i\in I_{\texttt{user}}$, $v_{i}$ is an emotion node, which encapsulates the fine-grained emotion state of the user. For agent turns $i\in I_{\texttt{agent}}$, $v_{i}$ is a system strategy node, that represents the specific conversational strategy implemented by the agent. Lastly, we introduce the dummy node $v_{N}$ as a placeholder for the target utterance, aggregating information from the two other node types and their interactions.
Edge Types: The edges in our heterogeneous graph fulfill dual roles: they model the discourse dependencies between dialogue turns and facilitate information aggregation towards the dummy node. Discourse dependencies correspond to edges between and within emotion and system strategy nodes: we follow Asher et al. (2016)’s definition (which include categories such as Comment and Elaboration) and pre-train a discourse parser as proposed by Chi and Rudnicky (2022) on the multi-party discourse dataset STAC (Asher et al., 2016). Details regarding the training procedure can be found in Appendix G. We note $\mathcal{R}_{\text{Discourse}}$ the set of possible dependencies; then, $\forall(i,j)$ such that $1\leq i,j\leq N-1,\langle v_{i},v_{j}\rangle\in\mathcal{R}_{\text{Discourse}}$. We give more details on these dependencies in Appendix A.2. The remaining edges are aggregating information from system strategy nodes, which we call self-reference: $\forall i\in I_{\texttt{agent}},\langle v_{i},v_{N}\rangle=r_{\text{self}}$; and from user emotion nodes, which we call inter-reference: $\forall i\in I_{\texttt{user}},\langle v_{i},v_{N}\rangle=r_{\text{inter}}$. We note $\mathcal{R}=\mathcal{R}_{\text{Discourse}}\cup\{r_{\text{self}},r_{\text{inter}}\}$ the set of edge types.
@@ 4.2.1 User Emotion Node Embedding: Mixed Emotion Method
Unlike previous work that relies on commonsense knowledge, we propose leveraging a pre-trained ERC model to predict emotion distributions from user utterances. We then utilize the knowledge contained in these distributions to create embeddings for fine-grained user states using a mixed-prototype approach.
@@
For DailyDialog, we categorize emotions into seven groups: Ekman’s six basic emotions plus Neutral, collectively referred to as $\mathcal{E}$. Additional statistical details about DailyDialog, along with our motivation for selecting it, are available in Appendix A.2. Detailed information regarding the implementation and training hyperparameters can be found in Appendix G.
@@
To utilize the information in the emotion label distribution more effectively, we employ a learnable temperature parameter $\tau$; details on the impact of the initialization of $\tau$ can be found in Appendix I.
@@ 4.3 Next Dialogue Strategy Prediction
Model ESConv AnnoMI
$\mathcal{M}$-F1$\uparrow$ $\mathcal{W}$$\uparrow \mathcal{B}$$\downarrow \mathcal{M}$-F1$\uparrow$ $\mathcal{W}$$\uparrow \mathcal{B}$$\downarrow$
Prompting LLMs LLaMA3-70B (Meta, 2024) 15.36 18.45 1.03 8.38 9.01 1.24
+ 2 shot 17.70 21.47 1.29 9.52 9.13 1.11
+ ERC 15.70 19.32 1.12 8.28 9.02 1.30
ChatGPT (OpenAI, 2023) 18.14 20.27 0.88 20.31 18.12 1.21
+ 2 shot 16.55 20.01 0.73 15.29 14.22 1.39
+ ERC 16.50 18.79 0.77 16.17 15.60 0.89
Finetuning LLMs RoBERTa Liu et al. (2019) 25.04 27.94 0.68 22.26 27.25 0.64
BART (Lewis et al., 2020) 25.66 29.08 0.64 22.94 29.68 1.07
LLaMA3-8B (Meta, 2024) 25.91 29.82 0.83 23.77 29.98 0.81
Specialized Models MISC (Tu et al., 2022) 20.91 24.93 0.89 - - -
MultiESC (Cheng et al., 2022) 25.73 29.31 0.61 - - -
KEMI (Deng et al., 2023) 24.69 26.80 0.86 - - -
TransESC (Zhao et al., 2023) 26.28 31.33 0.73 - - -
EmoDynamiX ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$
Table 1: Experimental results on two ESC datasets. The best results are bolded and the second best are underlined. † indicates statistically significant improvement ($p<0.05$). Since MultiESC has adopted a new set of labels, we merge the updated ones with the original annotations to ensure a fair comparison. Due to the unavailability of TransESC’s data preprocessing pipeline, we report the reproduced results based on their released train/dev/test split.
Model ESConv AnnoMI
$\mathcal{M}$-F1$\uparrow$ $\mathcal{W}$$\uparrow \mathcal{B}$$\downarrow \mathcal{M}$-F1$\uparrow$ $\mathcal{W}$$\uparrow \mathcal{B}$$\downarrow$
EmoDynamiX 27.70 32.71 0.45 27.92 35.33 0.50
w/o Graph Learning 25.72↓1.98 29.31↓3.40 0.78↑0.33 26.95↓0.97 29.46↓5.87 0.73↑0.23
w/o Mixed Emotion 25.90↓1.80 29.45↓3.26 0.66↑0.21 24.71↓3.21 30.25↓5.08 0.70↑0.20
w/o Discourse Parser 26.64↓1.06 30.12↓2.59 0.59↑0.14 27.04↓0.88 31.59↓3.74 0.60↑0.10
w/o Dummy Node 25.46↓2.24 29.80↓2.91 0.73↑0.28 24.73↓3.19 29.00↓6.33 0.72↑0.22
Table 2: Evaluation results of ablation study.